# Omegance: A Single Parameter for Various Granularities in Diffusion-Based Synthesis

### v1

@@ 3.2.2 Omega Schedule
Figure 5: Global effects of Omegance in fixing visual artifacts and improving realism. (Zoom-in for best view)
Figure 6: Temporal effects of schedule-based Omegance. Four quadrants in (a) are the same as those in Fig. 4. EXP1: More complex layout, slightly more fine detail. EXP2: More complex layout, less fine detail. COS1: Slightly less complex layout, less fine detail. COS2: Less complex layout, more fine detail. (Zoom-in for best view)
Figure 7: Spatial effects of mask-based Omegance in SDEdit results. (Zoom-in for best view)
@@ 4 Experiments
We examine the effectiveness of Omegance across various generative models and applications, including Stable Diffusion XL (SDXL) [16], RealVisXL-V5.0 [21], Stable Diffusion 3 (SD3) [5], FLUX [13], FreeU [22], SDEdit [15], ContolNet [28], ReNoise [7], SDXL-Inpainting [16], Latte [14], and AnimateDiff [8]. Implementations of these methods are based on Hugging Face’s Diffusers repository [^1] https://github.com/huggingface/diffusers. Note that we exclude omega mask experiments in purely text-guided methods due to the lack of prior knowledge about the output composition.
Figure 8: Spatial effects of mask-based Omegance in ControlNet results. (Zoom-in for best view)
Figure 9: Spatial effects of mask-based Omegance in ReNoise inversion results. (Zoom-in for best view)
Figure 10: Effects of Omegance in image inpainting task. (Zoom-in for best view)
@@ 4.1 Text-to-Image Generation
In addition to providing granularity control, Omegance also occasionally enhances the generated outputs, as shown in Fig. 5. In lower quality models, like SDXL [16], Omegance’s detail suppression effectively addresses artifacts in human body parts, particularly in intricate areas like fingers and arms. Meanwhile, For high-quality models like FLUX [13] which tend to produce over-smoothed results, Omegance’s detail enhancement improves realism by restoring fine-grained textures and intricate details.
We conducted a two-part user study with 102 participants to evaluate Omegance’s effectiveness in granularity control and impact on output quality (see supp. for details). In Part 1, participants are asked to rank three images with and without Omegance based on their granularity. The average rank accuracy reflects the effectiveness of Omegance in granularity control. In Part 2, participants select the higher-quality result from pairs generated with and without Omegance, and we report the percentage of votes favoring Omegance or insisting equal quality. The results in Tab. 1 demonstrate that Omegance achieves effective granularity control without degrading the base model’s quality.
Average Rank Accuracy Output Quality
Omegance 93.94% 81.98%
Table 1: Results of user study on granularity control effectiveness and output quality of Omegance.
@@
Omega Schedule. We show two distinct omega schedules in Fig. 4 and their effects in Fig. 1(c). To further demonstrate the effectiveness of the omega schedule in controlling the granularity of the output layout and fine detail simultaneously, we illustrate more continuous schedules in Fig. 6. In the given case, layout complexity is generally reflected by the composition of the chalet and the fine detail richness by the festival decorations and footprints in the snow.
Figure 11: Effects of Omegance in Text-to-Video results. (Zoom-in for best view)
@@ 4.2 Image-to-Image Generation
In image-to-image tasks, prior knowledge of image composition from reference inputs or structural guidance enables the effective use of omega masks to apply Omegance selectively. By assigning specific $\omega$ values to targeted regions, we achieve precise control over texture richness and smoothness, enhancing details in some areas while simplifying others. It is worth noticing that although unselected regions are not explicitly constrained, they experience minimal changes to maintain input consistency, demonstrating the precise region-based control of our omega mask. In all mask notations throughout this paper, red indicates detail enhancement, and blue represents detail suppression.
SDEdit. For image-to-image editing using SDEdit [15], we use off-the-shelf semantic segmentation tool SAM2 [18] to generate the segmentation masks for the input images, enabling the application of omega masks to specific segmentation regions. Results are demonstrated in Fig. 7.
ControlNet. Results of applying omega mask in ControlNet [28] are shown in Fig. 8. With ControlNet control signals, we can generate default masks that specify the regions of interest. For pose signal, which infers the position and pose of the main character, we apply dilated convolution on the skeleton to obtain a default character mask. For depth signal that conveys foreground and background information, we can use continuous depth values to create smooth, depth-wise transitions in the mask. Alternatively, it is always feasible to generate custom masks from user-provided strokes, allowing for more flexible and intuitive control over detail granularity. We use custom masks for canny signals.
Real Image Editing. Omegance can also be applied spatially in real-image inversion tasks to achieve granularity editing of specific objects. The masks can be obtained from either segmentation tools or user-provided strokes. The results using ReNoise inversion [7] are presented in Fig. 9.
Image Inpainting. We further generalize the use of Omegance to image inpainting task. Figure 10 shows the results of performing image inpainting using SDXL [16].
@@ 4.3 Text-to-Video Generation
Omegance’s granularity control ability can also generalize to text-to-video tasks. Experiments are conducted using Latte [14] and AnimateDiff [8]. In Fig. 11(a), “$\omega$ increasing” (detail suppression) leads to a less complex background and smoother texture, while “$\omega$ decreasing” (detail enhancement) corresponds to a more complex background and sharper texture. Additionally, current text-to-video generation techniques often suffer from visual artifacts in the output, as illustrated in the first row of Fig. 11(b), where the panda’s guitar is distorted on the left, and an unintended flow effect appears in the sky on the right. Applying Omegance can effectively address the above artifacts through its detail enhancement and suppression effects, respectively, as demonstrated in the second row of Fig. 11(b).
@@ 5 Conclusion and Limitation
We introduced Omegance, a simple yet effective single-parameter technique for controlling granularity in diffusion model outputs, enabling fine-grained spatial and temporal adjustments to layout complexity and texture richness. The method is training-free, architecture-agnostic, and integrates seamlessly with various diffusion-based tasks. Extensive experiments demonstrate Omegance’s ability in controlling level of detail in text-to-image, image-to-image, and text-to-video generation results. While Omegance excels at nuanced granularity manipulation and can occasionally correct artifacts or enhance realism, it does not inherently improve the generation quality of the base model, which remains a limitation. Nevertheless, we believe this work is valuable for advancing controllable and user-driven content generation, expanding the practical applications of diffusion-based synthesis in real life.
Supplementary Material
This supplemental material mainly contains:
  • Representations of $\delta_{t}$ and $\zeta_{t}$ in Equ. 2 from different schedulers in Sec. A.
  • Step-by-step derivations of the modified SNR in Sec. B.
  • Detail description of user study setup in Sec. C.
  • More discussions in Sec. D.
  • More implementation details in Sec. E.
  • More qualitative results in Sec. F.
Appendix A Representations of $\delta_{t}$ and $\zeta_{t}$
Equation 2 (recapped below) provides a general representation of one denoising step in the diffusion reverse process. We provide detailed formulas of how to convert different denoising schedulers to the general representation.
$z_{t-1}=\delta_{t}\cdot z_{t}+\underbrace{\zeta_{t}\cdot\epsilon_{\theta}(z_{t},t)}_{\text{``direction pointing to $z_{0}$''}}$
(1) DDIM scheduler [23]:
$\begin{split}z_{t-1}&=\sqrt{\alpha_{t-1}}\underbrace{\frac{z_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(z_{t})}{\sqrt{\alpha_{t}}}}_{\text{``predicted $z_{0}$''}}+\underbrace{\sqrt{1-\alpha_{t-1}}\cdot\epsilon_{\theta}^{(t)}(z_{t})}_{\text{``direction pointing to $z_{t}$''}}\\ &=\delta_{t}\cdot z_{t}+\zeta_{t}\cdot\epsilon_{\theta}(z_{t},t)\end{split}$
where
$\begin{split}\delta_{t}&=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\\ \zeta_{t}&=-\sqrt{\alpha_{t-1}}\cdot\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}+\sqrt{1-\alpha_{t-1}}\end{split}$
(2) Euler discrete scheduler [12]:
$\begin{split}z_{t-1}&=z_{t}+(\sigma_{t+1}-\hat{\sigma})\cdot\epsilon_{\theta}(z_{t},t)\\ &=\delta_{t}\cdot z_{t}+\zeta_{t}\cdot\epsilon_{\theta}(z_{t},t)\end{split}$
where
$\begin{split}\delta_{t}&=1\\ \zeta_{t}&=\sigma_{t+1}-\hat{\sigma}\end{split}$
(3) Flow-matching scheduler [5]:
$\begin{split}z_{t-1}&=z_{t}+dt\cdot v_{\theta}(z_{t},t)\\ &=\delta_{t}\cdot z_{t}+\zeta_{t}\cdot\epsilon_{\theta}(z_{t},t)\end{split}$
where
$\begin{split}\delta_{t}&=1\\ \zeta_{t}\cdot\epsilon_{\theta}(z_{t},t)&\approx dt\cdot v_{\theta}(z_{t},t)\end{split}$
Appendix B Modified SNR Derivations
In DDIM denoising process,
$\begin{split}z_{t-1}&=\sqrt{\alpha_{t-1}}\underbrace{\frac{z_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}}{\sqrt{\alpha_{t}}}}_{\text{``predicted $z_{0}$''}}+\underbrace{\sqrt{1-\alpha_{t-1}}\cdot\epsilon_{\theta}^{(t)}(z_{t})}_{\text{``direction pointing to $z_{t}$''}}\\ &\approx\sqrt{\alpha_{t-1}}\hat{z_{0}}+\sqrt{1-\alpha_{t-1}}\hat{\epsilon_{\theta}}\end{split}$
So the original SNR during the reverse process is still: $\mathrm{SNR(t-1)}=\frac{\alpha_{t-1}}{1-\alpha_{t-1}}$.
Equation 6 shows the modified SNR based on DDIM denoising equation. The step-by-step derivations are provided below:
$\begin{split}z_{t-1}^{\prime}&=\sqrt{\alpha_{t-1}}\frac{z_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\cdot\omega}{\sqrt{\alpha_{t}}}+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}\cdot\omega\\ \end{split}$
Substitute $z_{t}=\sqrt{\alpha_{t}}z_{0}+\sqrt{1-\alpha_{t}}\epsilon_{\theta}$ into the equation and isolate $z_{0}$ and $\epsilon_{\theta}$ terms, we get:
$\begin{split}z_{t-1}^{\prime}=Az_{0}+B\epsilon_{\theta}\end{split}$
where
$\begin{split}A&=\sqrt{\alpha_{t-1}}\\ B&=\frac{\sqrt{\alpha_{t-1}}\sqrt{1-\alpha_{t}}(1-\omega)+\sqrt{\alpha_{t}}\sqrt{1-\alpha_{t-1}}\omega}{\sqrt{\alpha_{t}}}\end{split}$
Therefore, the modified SNR is:
$\begin{split}\mathrm{SNR}^{\prime}&=\frac{A^{2}}{B^{2}}\\ &=\frac{\alpha_{t-1}}{\frac{(\sqrt{\alpha_{t-1}}\sqrt{1-\alpha_{t}}(1-\omega)+\sqrt{\alpha_{t}}\sqrt{1-\alpha_{t-1}}\omega)^{2}}{\alpha_{t}}}\\ &=\frac{\alpha_{t-1}}{\left(\frac{\sqrt{\alpha_{t-1}}\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}+\frac{\sqrt{\alpha_{t}}\sqrt{1-\alpha_{t-1}}-\sqrt{\alpha_{t-1}}\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\omega\right)^{2}}\\ &\propto\omega\end{split}$
Since $\alpha_{t}$ monotonically decreases in diffusion model, so $\alpha_{t-1}>\alpha_{t}$ and $1-\alpha_{t}>1-\alpha_{t-1}$. Therefore, $\sqrt{\alpha_{t}}\sqrt{1-\alpha_{t-1}}-\sqrt{\alpha_{t-1}}\sqrt{1-\alpha_{t}}$ is always negative and $\mathrm{SNR}^{\prime}$ increases as $\omega$ increases.
Figure 12: Change of omega (left) vs. Change of number of inference steps (right). Example 1. Orange box indicates default results when $\omega=1.0$ and the number of inference steps $=50$.
Appendix C User Study
To demonstrate the effectiveness of Omegance in controlling image granularity and preserving image quality, we design a two-part user study.
In Part 1, for each question, the participants are given three images, one generated by the base model, and the other two by the base model with Omegance. The participants are asked to select the rank of image granularity (from high to low) that best matches their inspection. Instructions about granularity are given: Granularity in the context of image generation refers to the level of detail and texture richness in a visual output. High granularity corresponds to intricate textures, complex patterns, and rich visual density, while low granularity is associated with smoother transitions, minimal detail, and simpler compositions. The motivation behind such design is that if the granularity rank can be correctly distinguished by the users, it proves the effectiveness of Omegance in controlling image granularity. The images used in Part 1 are generated by SDXL [16] and FLUX [13] using global Omegance. For quantitative evaluation, we consider each rank position separately and calculate the accuracy independently. The final average rank accuracy is the average accuracy of all rank positions.
In Part 2, our objective is to illustrate that Omegance does not harm the base model’s quality and can occasionally improve it by fixing artifacts and enhancing realism. We present two contents generated by the base model with and without Omegance and ask the users to select the one with better quality. The instructions on visual quality definition are given as: Visual quality refers to the overall perceptual appeal and coherence of a generated content, encompassing aspects such as sharpness, realism, fidelity to given prompt, and freedom from artifacts. We believe that by achieving over 50% votes in quality competition, it is sufficient to prove that Omegance can achieve at least equal quality outcomes as the base model. The results used in Part 2 are generated by SDXL [15], FLUX [13], RealVisXL-5.0 [21], AnimateDiff [8], and Latte [14]. We report the percentage of votes indicating our method achieves equal or higher quality over the base model as our quality evaluation.
Appendix D Discussion
In this section, we discuss several settings that achieve similar effects in granularity or motivate our design choices.
D.1 Change Inference Steps
Figure 13: Change of omega (left) vs. Change of number of inference steps (right). Example 2. The orange box indicates default results when $\omega=1.0$ and the number of inference steps $=50$.
An intuitive approach to influencing the granularity of generative outputs is to adjust the number of inference steps. However, as observed in our experiments, linearly increasing the number of steps does not provide consistent granularity control. Figure 12 and 13 demonstrate this using SDXL [16] with the Euler discrete scheduler [12], where changes in layout complexity and texture richness caused by increasing steps show irregular patterns compared to the consistent effects achieved with Omegance. Moreover, modifying the number of steps impacts the entire image globally, offering no capacity for region-specific adjustments. In contrast, Omegance enables localized control through omega masks and schedules. Omega masks allow region-specific granularity adjustments, enhancing or suppressing detail in specific regions while preserving unselected areas, and omega schedules can refine layout or texture granularity at different denoising stages. This flexibility is especially valuable for real-world design tasks, where different regions or elements require varying levels of detail. Besides, increasing steps adds computational overhead without guaranteeing significant improvements in granularity control, while Omegance provides granularity manipulation within a fixed inference schedule with no extra cost, maintaining computational efficiency and output fidelity.
In summary, adjusting inference steps is a coarse mechanism for granularity control, as it alters the overall denoising dynamics without offering fine-tuned control. Omegance, by operating directly within the existing denoising framework, delivers precise, user-driven adjustments without additional computational cost.
D.2 Change Latent Mean
Figure 14: Effects of latent mean changes on image RGB mean with analysis.
In Omegance, we carefully preserve the mean of $z_{t}$ while altering its variance. This design choice is motivated by our observation that changes in the latent mean lead to color shifts in the final decoded outputs. This phenomenon can be verified by directly modifying the mean of VAE-encoded latents in a latent diffusion model [19]. Given an original image in Fig. 14(a) as $x_{0}$, we encode the image to latent $z_{0}$ using the VAE used in SD3 [5] and modify the mean of $z_{0}$ directly. The results of increasing and decreasing latent mean are shown in Fig. 14(b) and (c), respectively. These results reveal that the green channel is particularly sensitive to mean changes, while the red channel is the least affected. Consequently, increasing the mean causes the green color to dominate, while decreasing it leads to a dominance of red tones, producing undesirable color shifts. To avoid such artifacts, we ensure mean preservation in the implementation of Omegance.
Appendix E More Implementation Details
In this section, the implementation details for producing the demo contents shown in the main paper are explained.
E.1 Prompts for Demo
We use Large Language Model to help generate prompts.
Figure 1(a): A serene lake at dawn, with crystal-clear water reflecting snow-capped mountains and a soft pink sky. In the foreground, delicate wildflowers bloom along the shoreline, while a small wooden rowboat drifts quietly. The smooth, mirror-like surface of the lake contrasts with the rich details of the flowers, boat, and distant mountains, all softly illuminated by the gentle morning light.
Figure 1(b): A wise wizard in a long, starry cloak stands in an enchanted forest clearing, with towering trees and vibrant mushrooms glowing softly. Around him, floating magical orbs and sparkling fireflies create a mystical ambiance, while wildflowers and ivy-covered stones add texture to the scene.
Figure 1(c): A bustling medieval marketplace filled with colorful tents and stalls, where merchants display spices, textiles, and jewelry. Cobblestone streets wind between the booths, and in the distance, a towering castle rises against the horizon.
Figure 3 SDXL (Left): A gentle healer dressed in flowing robes stands beside a calm forest pond, her hand extended over the water as soft, glowing light surrounds her. The pond’s smooth surface reflects her figure and the towering trees around, while small, vibrant wildflowers dot the mossy shore.
Figure 3 SDXL (Middle): A peaceful Japanese zen garden at dusk, with smooth raked sand patterns surrounding moss-covered stones and a gentle, flowing stream reflecting the warm glow of lantern light.
Figure 3 SDXL (Right): A futuristic space station interior with rounded, smooth walls and control panels filled with colorful buttons and screens. Astronauts in sleek spacesuits float gently in zero gravity, surrounded by floating tools and shimmering holographic displays showing star maps and distant planets.
Figure 3 SDXL + FreeU: A bustling riverside café painted in impressionist style, where figures in soft, muted tones gather under glowing lanterns. The river reflects the lights, creating smooth, rippling patterns as the colors blend seamlessly with the surrounding trees and buildings.
Figure 3 RealVisXL-5.0: A serene meadow at sunrise, with a vintage picnic blanket spread out under a large oak tree. Wildflowers in pastel colors bloom across the soft grass, and a wicker picnic basket filled with freshly baked bread and fruit adds to the idyllic, pastoral scene.
Figure 3 SD3: A celestial observatory with smooth, polished floors and a massive domed ceiling painted with constellations in deep blues and silvers. A large telescope stands at the center, and intricate star maps and charts are scattered around, bathed in the soft, ambient light of glowing stars.
Figure 3 FLUX: A cozy anime café on a rainy afternoon, where customers sit at tables enjoying warm drinks as raindrops patter against the large windows. Soft, warm lighting gives the café a welcoming feel, with framed pictures, books, and cushions scattered around, adding rich detail to the snug interior.
Figure 5: A boy is playing Pokemon.
Figure 6: A quaint alpine village market in winter, with stalls selling handmade crafts, baked goods, and hot drinks, nestled between snow-covered wooden chalets. Pine trees and mountain peaks frame the background, while a soft snowfall adds charm to the festive, bustling atmosphere.
Figure 7 (Top): A creamy white marble table sprinkled with tiny violet and daisy blooms supports a classic white plate with thin, golden edges. The chocolate cake, layered with dark chocolate curls and topped with vibrant berries, sits beautifully on the plate, creating a striking contrast between the lush texture of the cake and the refined elegance of the setting.
Figure 7 (Bottom): A cool fox wearing sunglasses leaning on a rusted water pipe.
Figure 8 SDXL Pose (Left): Darth vander dancing in a desert, high quality.
Figure 8 SDXL Pose (Right): An elven warrior princess stands in a lush forest glade, her armor decorated with delicate leaf patterns and shining gemstones. Her cape is embroidered with nature motifs, and her ornate bow and quiver add a touch of elegance, creating a striking blend of nature and nobility.
Figure 8 SDXL Depth (Left): A topiary plant decorated by flowers.
Figure 8 SDXL Depth (Right): A sleek, futuristic robot with a polished metallic body and glowing blue eyes, standing upright with articulated limbs and fine details in its joints and circuitry. Soft lights reflect off its surface, highlighting its advanced design and smooth, streamlined form.
Figure 8 SDXL Canny (Left): Aerial view, a futuristic research complex in a bright foggy jungle, hard lighting.
Figure 8 SDXL Canny (Right): A stunning piece of jewelry featuring a delicate, silver pendant encrusted with sparkling diamonds and a large, brilliant-cut sapphire at its center, suspended from a fine, intricately detailed chain that glimmers in soft light.
Figure 8 FLUX Canny (Right): A sparkling crystal mansion stands in the middle of a blooming meadow, its walls refracting sunlight into brilliant rainbows. Delicate cherry blossom trees line the cobblestone path leading to the ornate glass doors, while butterflies flutter in the gentle breeze.
Figure 8 FLUX Canny (Left): A whimsical, pastel-colored cottage with a gently sloping roof and rounded windows sits amidst a vibrant garden of oversized flowers. A cobblestone path leads to the arched wooden door, and sparkling lights hang from the eaves, glowing softly under the golden afternoon sun.
Figure 9 (Top): A glass vase filled with flowers.
Figure 9 (Bottom): A kitten in a basket.
Figure 10: A loosely arranged bouquet of wildflowers in a glass jar sits in the center of the bench, with petals and leaves spilling over the edge. The soft textures of the flowers balance the structure of the jar and bench for a touch of elegance.
E.2 Negative Prompt
We use negative prompts where applicable to help improve generation quality: “distorted lines, warped shapes, uneven grid patterns, irregular geometry, misaligned symmetry, low quality, bad quality”.
E.3 Omega Rescale
Figure 15: Visualization of omega rescale function.
As mentioned in Sec. 3.2, we rescale omega to allow finer-grained control within (,)($\infty$,$\infty$)( - ∞ , ∞ ) input range by:
$\omega=\mathcal{R}(\varpi)=L+\frac{U-L}{1+e^{-k\cdot\varpi}}$ (10)
The default rescaling function for SDXL [16] model is visualized in Fig. 15 with $k=0.1,L=0.95,U=1.05$. In this case, $\varpi$ within $[-10.0,10.0]$ would present visible effects. In SD3 [5] and FLUX [13], the range of $\omega$ should be larger around $[0.8,1.2]$.
Appendix F More Qualitative Results
In this section, we show more examples of Omegance in SDXL [16], SDXL+FreeU [22], RealVisXL-V5.0 [21], SD3 [5], FLUX [13], and ControlNet [28]. Control signals are indicated in the top left corner of the original results. Omega masks used are indicated in the bottom left corner of the corresponding Omegance-edited results. Red for detail enhancement, blue for detail suppression.
Figure 16: More global effects of Omegance in SDXL.
Figure 17: More global effects of Omegance in SDXL+FreeU and RealVisXL-V5.0.
Figure 18: More global effects of Omegance in SD3 and FLUX.
Figure 19: More temporal effects of schedule-based Omegance follow schedules defined in Fig. 6.
Figure 20: More spatial effects of mask-based Omegance.
Figure 21: More spatial effects of mask-based Omegance using ControlNet with canny signal.

### v2

@@ 3.2.2 Omega Schedule
Figure 5: The effects of global Omegance in fixing visual artifacts and improving realism.
Figure 6: Temporal effects of schedule-based Omegance. Four quadrants in (a) are the same as those in Fig. 4. Four examples are illustrated: EXP1: More complex layout, slightly more fine detail. EXP2: More complex layout, less fine detail. COS1: Slightly less complex layout, less fine detail. COS2: Less complex layout, more fine detail. (See implementation details in supp. Sec.D)
@@ 4 Experiments
We examine the effectiveness of Omegance across various generative models and applications, including Stable Diffusion XL (SDXL) [30], RealVisXL-V5.0 [37], Stable Diffusion 3 (SD3) [8], FLUX [22], FreeU [38], SDEdit [26], ControlNet [48], ReNoise [10], SDXL-Inpainting [30], Latte [25], and AnimateDiff [11]. Implementations of these methods are based on Huggingface’s Diffusers repository [^2] https://github.com/huggingface/diffusers. More implementation details are in supp. Sec.J.
@@ 4.1 Text-to-Image Generation
Table 1: Quantitative comparisons of image quality, text-image alignment, and aesthetics with previous works. Different Omegance settings are highlighted by a blue background. Bold and underline indicate best and second best results, respectively.
[…]
SDXL [30] 162.18 13.23 32.88 4.68 0.1468
+ FreeU [38] 167.22 12.25 31.76 4.64 0.0967
+ Cosine Sch. [28] 182.06 11.38 30.78 2.88 0.0376
+ Rescaled Sch. [24] 163.29 10.88 28.80 3.25 0.0295
+ $\varpi(6.0)$ 157.47 13.82 32.70 4.64 0.1149
+ $\varpi(-6.0)$ 170.52 13.01 32.81 4.67 0.1601
+ EXP1 173.49 12.67 32.70 4.64 0.1578
+ COS1 159.87 13.25 32.64 4.60 0.0962
Table 2: High-Frequency Energy (HFE) of different Omegance settings and their changes w.r.t. SDXL baseline in brackets.
[…]
SDXL [30] 1.0 1159.4 (0)
+ $\varpi(6.0)$ 0.8124 955.2 (-204.2)
+ $\varpi(-6.0)$ 0.7940 1680.1 (+520.7)
+ EXP1 0.7087 2272.5 (+1113.1)
+ EXP2 0.6926 1365.2 (+205.8)
+ COS1 0.8183 1004.7 (-154.7)
+ COS2 0.7311 612.8 (-546.6)
Table 3: Win rate of Omegance compared to baselines in granularity control effectiveness and output quality.
Average Rank Accuracy Output Quality
Omegance 93.94% 81.38%
Figure 7: Spatial effects of mask-based Omegance in ControlNet results. Omegance enables spatially controlled granularity adjustments while preserving untouched areas. Mask annotation: red indicates detail enhancement, blue represents detail suppression, and white denotes unchanged regions.
@@
In addition to providing granularity control, Omegance also occasionally enhances the generated outputs, as shown in Fig. 5. In lower-quality models, like SDXL [30], Omegance’s detail suppression effectively addresses artifacts in human body parts, particularly in intricate areas like fingers and arms. Meanwhile, for high-quality models like FLUX [22], which tend to produce over-smoothed results, Omegance’s detail enhancement improves realism by restoring fine-grained textures and intricate details.
@@
Omega Schedule. We show two discrete omega schedules in Fig. 4 and their effects in Fig. 1(c). To further demonstrate the effectiveness of omega schedule in controlling the granularity of the output layout and fine detail simultaneously, we illustrate more continuous schedules in Fig. 6. In the given case, layout complexity is generally reflected by the composition of the chalet and the fine detail richness by the festival decorations and footprints in the snow. Note that these omega schedules are merely illustrative and not exhaustive. One can design own omega schedules by following the effects demonstrated in Fig. 4.
Quantitative Results. To assess the effectiveness of Omegance, we conducted experiments using 1,000 randomly sampled prompts from DiffusionDB [43], with SDXL as the base model. To ensure a comprehensive evaluation, we report Fréchet Inception Distance (FID) [13] and Inception Score (IS) [35] for image quality, CLIP score [31] for text-image alignment, and Q-Align [44] and PickScore [20] for aesthetics. As presented in Tab. 2, Omegance (highlighted in blue) outperforms structure-modification methods [38] and scheduler-based approaches [28, 24] across all key dimensions. Generally, detail suppression (e.g., $\varpi=6.0$ and COS1) enhances image quality, as indicated by lower FID and higher IS scores compared to the base SDXL model. Conversely, detail enhancement (e.g., $\varpi=-6.0$ and EXP1) maintains CLIP and Q-Align scores on par with SDXL, while significantly improving aesthetic appeal, as reflected in higher PickScore values.
To validate that Omegance preserves overall image composition, we report structural similarity (SSIM) [42] in Tab. 2, confirming that it minimally alters the global layout, a finding also supported by qualitative results. Analyzing high-frequency components is integral to various image processing applications. Here, we perform frequency-domain analysis to quantify fine-detail changes relative to the base model (formulation in supp. Sec.G). A higher mean high-frequency energy (HFE) ($+$) indicates enhanced details, while a lower HFE (--) reflects detail suppression. As shown in Tab. 2, Omegance’s global effects (2nd and 3rd rows) align with our analysis in Sec. 3.2. In addition, the HFE values in rows 4-7 show that the metric conforms closely to the different omega schedules we tested in Fig. 6. If consistent effect is applied to both early and late inference stages (e.g., EXP1 and COS1), HFE changes align with the effect accordingly. However, when early and late-stage effects differ (e.g., EXP2 and COS2), the resulting HFE changes depend on the severity and duration of each effect across the inference stages. This highlights the flexibility of designing custom omega schedules to selectively control detail granularity in layout structure and texture refinements, based on specific generation requirements.
Less Detail Original More Detail
[…]
Refer to caption Refer to caption Refer to caption
(a) Mochi: “Cinematic close-up shot of a sad woman riding a bus in the rain, cool blue tones, sad mood.”
[…]
Refer to caption Refer to caption Refer to caption
(b) Hunyuan: “A cute creature with snow leopard-like fur is walking in winter forest, 3D cartoon style render.”
Figure 8: Effects of Omegance in Text-to-Video results. The employed T2V models and prompts are shown below, with the detail variations on top. Omegance enables granularity control in generated videos while preserving temporal coherence. (Use Adobe PDF Reader for animated view.)
User Study. In addition to metric evaluation, we conducted a two-part user study with 101 participants to evaluate Omegance’s effectiveness in granularity control and impact on output quality (supp. Sec.H for details). In Part 1, participants are asked to rank three images with/without Omegance based on their granularity. Average rank accuracy reflects the effectiveness of Omegance in granularity control. In Part 2, participants select the higher-quality result from image pairs with/without Omegance, and we report the percentage of votes favoring Omegance or insisting on equal quality. The results in Tab. 3 demonstrate that Omegance achieves effective granularity control without degrading base model’s quality. Notably, 81.38% of users responded positively to Omegance results, with 67.62% favoring them and 13.76% finding them comparable.
@@ 4.2 Image-to-Image Generation
In image-to-image tasks, prior knowledge of image composition from reference inputs or structural guidance enables the effective use of omega masks to apply Omegance selectively. By assigning specific $\omega$ values to targeted regions, we achieve precise control over texture richness and smoothness, enhancing details in some areas while simplifying others. It is worth noticing that although unselected regions are not explicitly constrained, they experience minimal changes to maintain input consistency, demonstrating the precise region-based control of our omega mask.
ControlNet. Results of applying omega mask in ControlNet [48] are shown in Fig. 7. With ControlNet control signals, we can generate default masks that specify the regions of interest. For the pose signal, which infers the position and pose of the main character, we apply dilated convolution on the skeleton to obtain a default character mask. For a depth signal that conveys foreground and background information, we can use continuous depth values to create depth-aware masks. Alternatively, it is always feasible to generate custom masks from user-provided strokes, allowing for more flexible and intuitive control over detail granularity. We use custom masks for canny signals.
Other tasks. More results of SDEdit [26], ReNoise [10], and SDXL-Inpainting [30] are shown in supp. Sec.E.
@@ 4.3 Text-to-Video Generation
Omegance’s granularity control ability also generalizes to text-to-video applications, e.g., Mochi [41] and Hunyuan [21]. In Fig. 8, “Less Detail” ($\omega$ increasing) leads to a less complex background and smoother texture, which highlights the main character. On the contrary, “More Detail” ($\omega$ decreasing) corresponds to a more complex background and sharper texture, like raindrops on the window in (a) and snow on the ground in (b), leading to more realistic visual results. More text-to-video results of Latte [25] and AnimateDiff [11] are shown in supp. Sec.F.
@@ 5 Conclusion and Limitation
We introduced Omegance, a simple yet effective single-parameter technique for controlling granularity in diffusion model outputs, enabling fine-grained spatial and temporal adjustments to layout complexity and texture richness. The method is training-free and architecture-agnostic and integrates seamlessly with various diffusion-based tasks. Extensive experiments demonstrate Omegance’s ability to control the level of detail in text-to-image, image-to-image, and text-to-video generation results. While Omegance excels at nuanced granularity manipulation and can occasionally correct artifacts or enhance realism, it does not inherently improve the generation quality of the base model, which remains a limitation. However, its simplicity is a strength: Unlike existing methods that require model retraining or complex architecture modifications, Omegance leverages a fundamental operation—noise scaling—in a previously unexplored way to enable effective control over generation granularity. By demonstrating its versatility across various tasks, we establish that even a simple modification, when properly applied, can lead to substantial improvements in controllability and user interaction with generative models. Nevertheless, we believe this work is valuable for advancing controllable and user-driven content generation, expanding the practical applications of diffusion-based synthesis in real life.
Acknowledgement. This study is supported under the RIE2020 Industry Alignment Fund Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).